Actions
Meeting 1-3-2017¶
Date: 1st Mar 2017
Time: 15:30 UK time
Attendance:
Apologies: Paul K
Agenda and Minutes¶
Prev Actions:
Action: Ed, Spare Boards still need temp on RH boards and bias calib on LH boards. Will do it in a week in January.
New Actions:
Data Taking Summary MAU, EO, PK, AD¶
http://www.hep.ph.ic.ac.uk/~mgeorge/Files/2016MICEData/
Dead channels DS¶
Efficiency DS Round table discussion,¶
Descoped cooling Tracker design - AD¶
Status of spares: EO CMW
- cold heads and boards
4 new absorbers have been ordered and will arrive for the shutdown and be installed on each cryostat.
Quotes have been requested for two new compressors.
2. Software: PK, AD, CHunt, CHeidt¶
Paul K send update in his absence that he is working on getting MAUS 2.8.3 working.
- Performance and Analysis
- Validation script
- Geometry,
- Calibrations, Configurations
- Monte Carlo
- Pattern Recognition
- Kalman
- Online Displays
AoB¶
Phone Details: International Toll-Free Dial-in numbers¶
- Switzerland: 0800705351
- UK: 08004960576
- US: 1-866-740-1260
- Other numbers
- Access Code: 5673390
Updated by George, Melissa over 6 years ago ยท 3 revisions